A planet orbiting a distant star has a radius 3.24 Γ— 106 π‘š. The escape speed for an object launched from this planet’s surface is 7.65 Γ— 103 π‘š/𝑠. What is the acceleration due to gravity at the surface of the planet?


1
Expert's answer
2021-12-20T17:57:21-0500

Let's first find the mass of the planet:


"v=\sqrt{\dfrac{2GM}{R}},""v^2=\dfrac{2GM}{R},""M=\dfrac{v^2R}{2G},""M=\dfrac{(7.65\times10^3\ \dfrac{m}{s})^2\times3.24\times10^6\ m}{2\times6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}}=1.42\times10^{24}\ kg."

Finally, we can find the acceleration due to gravity at the surface of the planet:


"g=\dfrac{GM}{R^2},""g=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times1.42\times10^{24}\ kg}{(3.24\times10^6\ m)^2}=9.02\ \dfrac{m}{s^2}."